OpenStudy (anonymous):

14a + 10a = 72 what is the answer

ganeshie8 (ganeshie8):

you can add "like terms"

ganeshie8 (ganeshie8):

"14a" and "10a" are like terms, so u can add them

ganeshie8 (ganeshie8):

whats 14 + 10 ?

OpenStudy (anonymous):

24

ganeshie8 (ganeshie8):

yes ! so ur left hand side becomes : 14a + 10a = 24a

ganeshie8 (ganeshie8):

ur equation becomes : 24a = 72

ganeshie8 (ganeshie8):

ur goal is to solve "a"

ganeshie8 (ganeshie8):

any idea wat we can do to isolate "a" from below :- 24a = 72 ?

OpenStudy (anonymous):

the answer to your problem is prayer

OpenStudy (anonymous):

so what is the answer

ganeshie8 (ganeshie8):

isolate "a", u wud get the answer

ganeshie8 (ganeshie8):

24a = 72

ganeshie8 (ganeshie8):

to isolate "a", simply divide 24 both sides

OpenStudy (anonymous):

is it a=3

ganeshie8 (ganeshie8):

24a/24 = 72/24 a = 3

ganeshie8 (ganeshie8):

Awesome, yes ! good job :)

OpenStudy (anonymous):

what is 2(2x - 4) + x = 7

ganeshie8 (ganeshie8):

first work the parenthesis : multiply 2 wid the contents of parenthesis

ganeshie8 (ganeshie8):

2(2x - 4) + x = 7 2(2x) - 2(4) + x = 7 4x - 8 + x = 7

ganeshie8 (ganeshie8):

fine so far ? :)

OpenStudy (anonymous):

x=?

ganeshie8 (ganeshie8):

next, group like terms

ganeshie8 (ganeshie8):

"4x" and "x" are like terms

ganeshie8 (ganeshie8):

2(2x - 4) + x = 7 2(2x) - 2(4) + x = 7 4x - 8 + x = 7 4x + x - 8 = 7 5x - 8 = 7

ganeshie8 (ganeshie8):

see if u can do the rest :)

ganeshie8 (ganeshie8):

u just need to isolate "x" by itself

OpenStudy (anonymous):

I can't do the rest

ganeshie8 (ganeshie8):

its okay, add 8 both sides and can u tell me wat u get ?

ganeshie8 (ganeshie8):

5x - 8 = 7 +8 +8

ganeshie8 (ganeshie8):

you wud get : 5x = 15

ganeshie8 (ganeshie8):

next, divide 5 both sides : 5x = 15 /5 /5

ganeshie8 (ganeshie8):

watever u get is the final answer :)

OpenStudy (anonymous):

4

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

whats 15/5 ?

OpenStudy (anonymous):

3

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

5x = 15 /5 /5 x = 3

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

np.. u wlc :)

OpenStudy (anonymous):

could you show work

OpenStudy (anonymous):

all together

ganeshie8 (ganeshie8):

2(2x - 4) + x = 7 2(2x) - 2(4) + x = 7 4x - 8 + x = 7 4x + x - 8 = 7 5x - 8 = 7 5x = 15 x = 3

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

u wlc :)